Find the three angles of a triangle if the second angle is twenty more than the first angle, and the third angle is twice the frist angle.
Answer by Fombitz(32388)

Let’s name the angles A, B, and C.
Then you know that B = A + 20 and C = 2(A).
Finally since it’s a triangle, the sum of all of the angles equals 180 or A + B + C = 180.
Substitute the values of B and C from above and you get :
A + B + C = 180
A + (A+20) + 2A = 180
4A + 20 = 180
4A + 20 – 20 = 180 – 20
4A = 160
A = 40
Substitute back into your equations for B and C.
B = A + 20
B = 40 + 20
B = 60
C = 2A
C = 2(40)
C = 80
+Final+Answers
A = 40
B = 60
C = 80